Muon Spectrometer White Board

Muon ATLAS Control Room 71363 & 71365
Muon Satellite Control Room 70948
Muon Run Coordinators: Stephanie Zimmermann (163050) and Enrico Pasqualucci (162197)

On-Call Experts

Primary On-call Experts -- your point of contact 24/7 -- Shifters, please stick to calling the numbers listed here and not other experts directly, except when instructed to do so by the primary on-call
SubDetector System On-call Number Period
Muon System run coordination/shift phone (**)   16-0226 permanent
MDT DAQ on-call Preema Pais 16-1939 Mar 28 - Apr 3
CSC/MDT DCS on-call see OTP 16-0376 permanent
      (16-3050) Backup
      (16-2858) CSC Backup
MDT Gas on-call Jianbei Liu 16-2532 Mar. 15 - Apr. 9
CSC DAQ on-call Kalliopi Iordanidou 0767951007
RPC LVL1/DAQ on-call   16-1853
RPC DCS/Detector on-call   16-0664
TGC General On-Call ... 16-1905  

(**) should there be no answer on the muon shift phone number, which has happened a few times due to problems with the network, try to call one of the muon run coordinators directly: Stephanie (163050), Enrico (162197), if this fails as well try Ludovico (163664)

Experts for specific items -- normally not to be called directly by the muon shifter

SubDetector System On-call Number Period
MDT GNAM/Online Monitoring Daniela Salvatore 16-2113  
    Gabriella Gaudio 16-2110  
TGC Gas Giora 16-0507  
TGC DCS Enrique 16-2530  
TGC DCS Shikma 16-5842  
TGC DCS Shlomit 16-4597  
TGC DAQ   16-3614  
TGC ROD Lorne / Daniel 16-0501 / 16-2431  

ATLAS Run plan

Additional Muon Plans and Muon Run Coordinator's notes

Note: This field should any special runs/tasks to be taken care of by the shifters. Please remove them when done.

  • TGC Recovery: As of March 26 Automatic Recovery of abandoned SSW (for RODs was in place already since a while) is active, please report problems.

LHC Scrubbing Run (April 1 to April ??, expect around 10 days)

  • CSC and TGC HV is off. .
  • MDT and RPC HV is at standard STANDBY voltage.
  • Due to HV off, CSC overall state in the DCS FSM appears as NOT_READY/OK. Check there are no other problems (LV, ...)
  • While HV is off, TGC overall state in the DCS FSM appears as READY, HV A and HV C FSM sub-trees have been excluded and locked out from the FSM.
  • During injection to verify muons are in safe state, relevant information is the safe for beam flag and muon injection permit, if the shift leader is worried on the TGC READY in the FSM tell him under this special conditions this is normal.
  • There will be no stable beams, thus no muon calibration stream, during the scrubbing run.

Shift Message Board

Shifters and Experts, when posting a comment please use a bullet (on a new line, type three spaces and then a *). Please add a date and your name so that we are aware that the information is recent

The Known Problems part in this section can be updated by everybody.

MDT Known Problems and Issues
  • 2011 - Missing MDT EE chambers: EE chambers other than EEL sector 5 A+C, EEL sector 13 A+C, EEL sector 11 side A, and EEL sector 15 side C are not yet installed and thus excluded or absent from the DCS FSM and disabled from readout in the MDT Control panel.
  • 2011 - OHP: MDTGlobalView tab: These histograms will only be filled when the stable beams flag is set
  • 2011.04.02 - Stephanie - BIS6A08 Mezz 0x200 disabled in the configuration DB, caused persistent chamber being dropped from the run. BIS1C16 mezz 0x800 disabled in the configuration DB, caused chamber being dropped frequently and errors in L1 order. Note: Expected mezz mask settings in DCS are not updated yet, there will be a WARNING in the JTAG status due to this for the 2 chambers.

CSC Known Problems and Issues
  • 2011 - No HV for Layers AL05 Layer 2 and AL09 Layer 4: HV permanently OFF due to broken wires.
  • 2011 - No DCS temperature monitoring for sectors 11,12 side A, sectors disabled in the DCS FSM (Temperature subtree)

RPC Known Problems and Issues
  • 2015 - *Some holes can be visible in plots concerning SL, but there's nothing to worry about: trigger hits and hit maps are OK, but a further optimization has to be performed. As example, in the SL_output_Hit_Map or in the SL_TriggerSector_vs_Tower plots, the following sectors can be empty or can have very low statistics:
    • sector 2 side A
    • sector 12 side C
    • sector 12 side A
    • sector 14 side C
    • sector 14 side A

For a complete and updated map of the holes, the following link shows the towers (highlighted in red) not included in the run: https://twiki.cern.ch/twiki/bin/view/Atlas/L1BarrelMuonTriggerHardware

TGC Known Problems and Issues
You will see unnecessary FATAL messages on the shifter assistance. Name; tgc-rol-stopless-removed
Message; Stopless removal took place for DF.ROS.ROS-TGC-EC-00.ReadoutModule1[9]
Action; Call the TGC expert, now! A recovery should happen in at most 2 minutes. Check that no channel is missing from the resource status page, and make sure that all RCD applications are running. A TTCRestart might be needed to recover disabled components and/or absent RCD applications. Make an elog entry (TGC as system affected).

And following INFORMATION will also appear
Name; tgc-rol-recovered

Please ignore above two messages for a while, we will fix the SA rule.

Special Shifter Instructions and Points Of Attention

Expert Message Board

Experts, feel free to put items currently to be watched out for or under investigation here. Please include your name and a date.

MDT Issues to watch out for ....

  • 2011.02.16 - Stephanie - Temperature for BrCtrl 14 Crate 0 Boards slightly high, monitor, abnormal low flow reading for cooling BW side A lower part racks so far thought as read problem, not actual low flow. Temp alarm adjusted on March 2 to 42 degree. Update March 16: Temperature is gradually increasing, also for rack YM11EMXA turbine temp in CIC. Will be investigated in upcoming tech stop.

  • 2011.10.13 - Stefanos Leontsinis - MDT PS 48V Y0805S2 Generator1 FanStatus, this alarm appears frequently. To be investigated in the next technical stop. Please keep an eye on this.

Last runs in the ATLAS partition (under development)



Last page update 01 Dec 2020 10:26:40


-- Main.EnricoTassi - 02-Apr-2011

Edit | Attach | Watch | Print version | History: r4 < r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r4 - 2018-08-28 - KentaroKawade
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Sandbox All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback